OpenStudy (anonymous):

if you divide 2sin2x by 2 do you get sinx or sin2x?

zepdrix (zepdrix):

Like this? \[\large \frac{2\sin(2x)}{2}\] You only perform the division once, meaning ~ on one of the factors in our multiplication on top.\[\Large =\sin(2x)\] Think about this to reinforce that idea,\[\Large \frac{2\cdot2}{2} =2\]We don't divide each of the 2's in the numerator by the 2 in the denominator.

OpenStudy (anonymous):

yes. thank you

OpenStudy (anonymous):

Wait. can you help me finish the problem? Im confused

zepdrix (zepdrix):

What is the rest of the problem? :o

OpenStudy (anonymous):

okay it was 2sin2x+sqrt3=0 I subtracted the sqrt of 3 and got 2sin2x=-sqrt3 then I divided by 2 and got sin2x=-sqrt3/2 Now Im not sure what to do next

OpenStudy (anonymous):

I have to find the 4 solutions between 0 and 2pi

zepdrix (zepdrix):

What is the rest of the problem? :o

OpenStudy (anonymous):

Thats it. It said find the 4 solutions between 0 and 2pi

zepdrix (zepdrix):

AHHHHHHHHHHH my openstudy crashed again :( lost all that typing grrr

zepdrix (zepdrix):

\[\Large \sin(2x)=-\frac{\sqrt3}{2}\]This value corresponds to a special angle on the unit circle. This next step might be a little bit confusing since our angle in this case is not \(\large \theta\) but instead is \(\large 2x\). Do you remember which places on the unit circle will give us this value when we take the sine?

OpenStudy (anonymous):

Take the sin of -sqrt3/2?

OpenStudy (anonymous):

Do you just ignore the 2 before the x?

zepdrix (zepdrix):

Let's ignore it for now, \[\large \sin u=-\frac{\sqrt3}{2}\] If you had this expression, what special angles do we get?

OpenStudy (anonymous):

4pi/3 and 5pi/3

zepdrix (zepdrix):

Good, we would say our `angle` \(\large u=4\pi/3\) and \(\large u=5\pi/3\). But our angle in this problem is the thing inside of the sine function. So we end up with,\[\large 2x=4\pi/3 \qquad\text{and}\qquad 2x=5\pi/3\]

zepdrix (zepdrix):

Our next step would be to solve for x. Confused by the 2x thing at all? :O

OpenStudy (anonymous):

yes.. Im lost with that 2 there

OpenStudy (anonymous):

so wait. are 4pi/3 and 5pi/3 solutions at all?

zepdrix (zepdrix):

\[\large \sin(Angle)=-\frac{\sqrt3}{2}\]And you determined that,\[\Large Angle=4\pi/3 \qquad\text{and}\qquad Angle=5\pi/3\] See how our angle in this problem is 2x? \[\Large \sin(2x)=-\frac{\sqrt3}{2}\]Giving us, \[\Large 2x=4\pi/3 \qquad\text{and}\qquad 2x=5\pi/3\]

zepdrix (zepdrix):

No they are not solutions. We haven't determined solutions yet :o